### v1

# Phonon Induced Energy Relaxation in Quantum Critical Metals

Abstract
Metals at the brink of electronic quantum phase transitions display high-temperature superconductivity, competing orders, and unconventional charge transport, revealing strong departures from conventional Fermi liquid behavior. Investigation of these fascinating intertwined phenomena has been at the center of research across a variety of correlated materials over the past many decades. A ubiquitous experimental observation is the emergence of a universal timescale that governs electrical transport and momentum relaxation. In this work, we analyze an equally important theoretical question of how the energy contained in the electronic degrees of freedom near a quantum phase transition relaxes to the environment via their coupling to acoustic phonons. Assuming that the bottleneck for energy dissipation is controlled by the coupling between electronic degrees of freedom and acoustic phonons, we present a universal theory of the temperature dependence of the energy relaxation rate in a marginal Fermi liquid. We find that the energy relaxation rate exhibits a complex set of temperature-dependent crossovers controlled by emergent energy scales in the problem. We place these results in the context of recent measurements of the energy relaxation rate via non-linear optical spectroscopy in the normal state of hole-doped cuprates.
Introduction.- In recent years, the problem of strange metals has drawn significant attention due to their unconventional transport behavior [1, 2, 3]. An experimental hallmark of these correlated metallic systems is a linear-in-temperature ($T$) electrical resistivity governed by a so-called Planckian relaxation rate of order $k_{B}T/\hbar$ down to low temperatures —in stark contrast to an umklapp-dominated $T^{2}$-scaling predicted by Fermi liquid theory, or a $T^{5}$-Bloch-Grüneissen regime due to scattering off acoustic phonons [4, 5]. The electrical transport timescale reflects the rate at which momentum is lost from the electronic system to external degrees of freedom, such as the lattice or impurities. Given the scale-invariant nature of the Planckian scattering rate [6] and experimental evidence for electronic quantum criticality across these materials [7, 8, 9, 10], a vast majority of theoretical efforts have focused on a scenario involving a Fermi surface coupled to the long-wavelength fluctuations of a bosonic collective mode of electronic origin [11, 12, 13], e.g. an electronic nematic [14]. The strong quantum critical fluctuations destroy the long-lived electronic quasiparticles near the Fermi surface, giving rise to non-Fermi liquid behavior. However, this does not automatically yield a $T-$linear momentum-relaxation rate, which has remained one of the central mysteries in the field. A number of theoretical works over the past few years have suggested additional microscopic ingredients that could potentially resolve this mystery [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
Figure 1: (a) The bottleneck for energy relaxation from the electronic d.o.f. at local temperature $T_{e}$ to the phonons at local temperature $T_{p}$ is governed by $\Gamma_{ep}$, which is controlled by the processes illustrated in (b)-(d). The lowest-order Feynman diagrams for the phonon self-energy due to the couplings: (b) ${\mathcal{L}}_{ep}^{(1)}$ Eq. (4a), (c) ${\mathcal{L}}_{ep}^{(2)}$ Eq. (4b), and (d) ${\mathcal{L}}_{ep}^{(3)}$ Eq. (4c), respectively. Solid (dashed) lines denote renormalized electron (bosonic collective mode, $\varphi$) Green’s function.
In this work, we focus on a related but even more poorly understood question: How does a strongly correlated metal relax its energy to the environment? While it is unlikely that phonons by themselves are responsible for the anomalous low-temperature momentum-relaxation in the strange-metal in the vicinity of quantum criticality [30], they are centrally important for energy relaxation (Fig. 1a). Phonons certainly act as a bath for the combined system consisting of electrons coupled to the bosonic collective mode, but have not received much theoretical attention in the context of non-Fermi liquids. The classic theoretical treatment [31] of energy relaxation rates, $\Gamma_{E}$, in a three-dimensional Fermi liquid coupled to acoustic phonons predicts a crossover from $\Gamma_{E}\propto T^{3}$ at low temperatures to $\Gamma_{E}\propto 1/T$ at high temperatures. In general, the momentum relaxation rate and $\Gamma_{E}$ are not automatically identical. How this behavior is modified in a non-Fermi liquid without long-lived electronic quasiparticles is presently unclear. We note that the theoretical framework developed below can be easily adapted to correlated insulating states that host Fermi surfaces of emergent neutral fermions (e.g. spinons) coupled to dynamical gauge-fields [32, 33], assuming that these degrees of freedom (d.o.f.) also relax their energy via coupling to acoustic phonons [34].
Model of thermal relaxation.- We propose a minimal low-energy model consisting of three d.o.f.: electrons, a bosonic collective mode of electronic origin tuned to the vicinity of quantum criticality, and acoustic phonons (Fig. 1a). Our goal is to understand the temperature dependence of energy transfer from the combined electronic degrees of freedom to the phonons. The new ingredient relative to previous considerations is the presence of the bosonic collective mode, which we expect to couple to phonons via symmetry-allowed channels. The direct coupling between the bosonic collective mode and the phonons in the low-energy effective theory are associated with a distinct kinematic regime, compared to any usual local electron-phonon coupling. We investigate the effects of three distinct couplings: (i) a direct deformation potential type coupling between the electrons and acoustic phonons, and (ii) the two leading-order (i.e. a linear and quadratic) couplings between the bosonic collective mode and the acoustic phonons, each motivated by symmetry and microscopic considerations. In general, the electron-phonon coupling can lead to a modification of the electronic quantum critical point [35, 36, 37, 38], but these effects are restricted to a small fraction of the Brillouin-zone [38]. Since energy relaxation involves phonons of all momenta, we will ignore these modifications; we also assume that these interactions are weak compared to the dominant electronic interactions.
Before introducing the low-energy theory, we first describe a two-temperature model for thermal relaxation, which forms the basis for all the calculations. The electron-electron ($\Gamma_{ee}$) and the phonon-phonon ($\Gamma_{pp}$) energy relaxation rates are assumed to be much faster than the electron-phonon ($\Gamma_{ep}$) energy relaxation rate, respectively. The energy relaxation rate from the phonons to the environment ($\Gamma_{p-env}$) is assumed to be an even slower process, i.e. $\{\Gamma_{ee},\Gamma_{pp}\}\gg\Gamma_{ep}\gg\Gamma_{p-env}$. The energy equilibration proceeds as follows: First, the electrons and the phonons reach their individual equilibria over a short time scale set by $\Gamma_{ee}$ and $\Gamma_{pp}$, with two different temperatures $T_{e}$ and $T_{p}$, respectively.  [^1] For acoustic phonons at low temperatures, the assumption of $\Gamma_{pp}\gg\Gamma_{ep}$ may no longer be satisfied. This implies that phonons at different energies will equilibrate with the electrons at different speeds. If the electrons and phonons start with significantly different initial temperatures, the transient phonon distribution function will be far from equilibrium. However, if the electrons are only driven slightly out of equilibrium, we expect $T_{e}\approx T_{p}$. Then phonons at different energies still have approximately similar effective temperatures, and our theory treatment should still approximately remain valid. Second, the electrons and phonons equilibrate with each other via $\Gamma_{ep}$, reaching a temperature $T_{\text{f}}$. Finally, the system dissipates heat into the environment with relaxation rate $\Gamma_{p-env}$. Therefore, our theoretical treatment will assume that the electron sector and the phonon sector equilibrate rapidly to their respective (different) temperatures, and the bottleneck for energy-relaxation is limited by $\Gamma_{ep}$ (Fig. 1a). Assuming $T_{e}\approx T_{p}\approx T$, the relaxation process is described by,
$C_{e}\frac{\partial T_{e}}{\partial t}$ $=$ $-\kappa\left(T_{e}-T_{p}\right)\,,$ (1a)
$C_{p}\frac{\partial T_{p}}{\partial t}$ $=$ $\kappa\left(T_{e}-T_{p}\right)\,,$ (1b)
where $C_{e}$ and $C_{p}$ are heat capacities of the electrons and phonons, respectively, and $\kappa$ characterizes the energy flux between the electron and phonon sectors. We have ignored here the losses to the environment via $\Gamma_{p-env}$. The non-zero eigenvalue of Eq. (1) is denoted $-\Gamma_{E}$, which is precisely the energy relaxation rate of interest,
$\Gamma_{E}=\kappa\left(\frac{1}{C_{e}}+\frac{1}{C_{p}}\right)\,.$ (2)
Note that the parametrically different dependence of $C_{e},~{}C_{p}$ on temperature inevitably accounts for some of the crossovers in $\Gamma_{E}$. To disentangle the thermodynamic $T-$dependent contributions to $C_{e},~{}C_{p}$ from the intrinsic $T-$dependence of the energy flux ($\kappa$) from electrons to phonons, in the remainder of this manuscript we focus primarily on $\kappa$ due to the distinct couplings described above. We will return to the full $T-$dependence of $\Gamma_{E}$ when we place our theoretical results in the context of the recent experiments [39].
Low-energy theory.- Consider a low-energy Lagrangian for electrons in the vicinity of quantum criticality, coupled to acoustic phonons via ${\mathcal{L}}={\mathcal{L}}_{e}+{\mathcal{L}}_{p}+{\mathcal{L}}_{ep}$, where
${\mathcal{L}}_{e}$ $=$ $\int{\rm d}\tau\Big{[}\sum_{\bm{k}}c_{\bm{k}}^{\dagger}\left(\partial_{\tau}+\varepsilon_{\bm{k}}\right)c_{\bm{k}}$ (3a)
$+\sum_{\bm{q}}\varphi_{-\bm{q}}\left(\partial_{\tau}^{2}+v_{\varphi}^{2}\bm{q}^{2}+r\right)\varphi_{\bm{q}}$
$+g\sum_{\bm{k},\bm{q}}f_{\bm{k},\bm{{q}}}c_{\bm{k+q/2}}^{\dagger}c_{\bm{k-q/2}}\varphi_{\bm{q}}\Big{]}+...,\,$
${\mathcal{L}}_{p}$ $=$ $\int{\rm d}\tau\sum_{i}\sum_{{\left|\bm{q}\right|}<{\omega_{\rm D}}/c}X_{i,-\bm{q}}\left(\partial_{\tau}^{2}+c^{2}\bm{q}^{2}\right)X_{i,\bm{q}}\,.$ (3b)
Here $c_{\bm{k}}^{\dagger}$ creates a fermion with momentum $\bm{k}$ and energy $\varepsilon_{\bm{k}}$, where $\sum_{\bm{k}}\equiv\int{\rm d}^{3}\bm{k}/(2\pi)^{3}$, and $\varphi_{\bm{{q}}}=\varphi_{-\bm{q}}^{\dagger}$ is the field operator for a quantum critical boson that carries no center of mass momentum (e.g. Ising-nematic order [14]). The boson velocity $v_{\varphi}$ is comparable to the Fermi velocity $v_{F}=\partial_{\bm{k}}\varepsilon_{\bm{k}}|_{k_{F}}$, and $r$ is used to tune to the quantum critical point (QCP) located at $r=r_{c}$. The Yukawa coupling of strength $g$ and form-factor $f_{\bm{k},\bm{q}}$ (e.g. $B_{1g}$ irreducible representation for the Ising-nematic) couples the electrons to the bosonic collective mode . The field $X_{i}$ denotes the $i$-th component of lattice displacement, and $c$ denotes the speed of sound. Here we have adopted the Debye model for phonons, i.e. the phonon momentum is cut off by ${\omega_{\rm D}}/c$, where ${\omega_{\rm D}}$ is the Debye frequency. For simplicity we have assumed that the different phonon polarizations have the same velocity. The ‘...’ in Eq. 3a represent a variety of generic perturbations, that can include disorder. Starting from Eq. (3a), a number of field-theoretic approaches yield a dynamical exponent $z_{\varphi}=3$ for the Landau-damped boson, and a non-Fermi liquid with $z_{c}=3/2$ in the clean theory [11, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 19, 72, 73, 74, 75, 17, 76, 20]. Inspired by a large body of experimental work [13], our goal is to focus on a scenario where in the absence of the coupling to phonons, the electrons are described by marginal Fermi liquid (MFL)-like correlations [77] in the quantum critical region. A variety of theoretical routes [43, 78], including disorder [71, 70, 19, 18], lead to $z_{\varphi}=2$, yielding an imaginary time MFL self-energy for the electrons, $\Sigma_{c}(i\omega)\sim-\omega\ln(\omega)$. While the exact microscopic mechanism leading to the ubiquitous evidence for MFL correlations is presently unclear, the following analysis of energy relaxation due to coupling to the acoustic phonons will be agnostic to the origin of MFL itself, as long as the electron-phonon interactions are weak.
We note that inspired in part by the recent experiments in cuprates, as well as other quasi two-dimensional correlated metals which display strange metallicity, here we will assume that while the electronic degrees of freedom are decoupled in the $z$-direction (i.e. the direct electronic hopping between layers is ignored), the phonons are entirely three-dimensional. We shall now include the three distinct channels for electron-phonon couplings, ${\mathcal{L}}_{ep}={\mathcal{L}}_{ep}^{(1)}+{\mathcal{L}}_{ep}^{(2)}+{\mathcal{L}}_{ep}^{(3)}$, where
${\mathcal{L}}_{ep}^{(1)}$ $=$ $\int{\rm d}\tau\sum_{\bm{k},\bm{q}}M^{i}_{\bm{k},\bm{q}}c^{\dagger}_{\bm{k+q/2}}c_{\bm{k-q/2}}X_{i,\bm{q}}\,,$ (4a)
${\mathcal{L}}_{ep}^{(2)}$ $=$ $\int{\rm d}\tau\sum_{\bm{q}}N_{\bm{q}}^{i}X_{i,-\bm{q}}\varphi_{\bm{q}}\,,$ (4b)
${\mathcal{L}}_{ep}^{(3)}$ $=$ $\frac{1}{2}\int{\rm d}\tau\sum_{\bm{k},\bm{{q}}}L^{i}_{\bm{k},\bm{q}}\varphi_{-\bm{k-q/2}}\varphi_{\bm{k+q/2}}X_{i,\bm{q}}\,.$ (4c)
Here, ${\mathcal{L}}_{ep}^{(1)}$ represents the standard electron-phonon coupling, where the matrix element $M^{i}_{\bm{k},\bm{q}}$ originates from coupling the lattice strain $(\partial_{i}X_{j}+\partial_{j}X_{i})$ to the electron density, and it encodes the correct linear combination in the symmetry allowed channel. ${\mathcal{L}}_{ep}^{(2)}$ is a direct linear coupling between the strain and the bosonic collective mode; such a coupling is allowed when the strain and $\varphi$ transform in the same fashion under crystalline symmetries, e.g. the nematic order can couple to the $B_{1g}$-strain modes [35]. Finally, ${\mathcal{L}}_{ep}^{(3)}$ represents the inelastic scattering process between the phonon and the bosonic collective mode [36]. According to Goldstone’s theorem, the matrix elements $M^{i}_{\bm{k},\bm{q}},N_{\bm{q}}^{i}$ and $L^{i}_{\bm{k},\bm{q}}$ are expected to be linear in the phonon momentum $\bm{q}$. For the coupling to electrons $M^{i}_{\bm{k},\bm{q}}$, the energy flux is neither sensitive to the specific form of $M^{i}_{\bm{k},\bm{q}}$ beyond the fact that $M^{i}_{\bm{k},\bm{q}}\propto\bm{q}$, nor to the quasi-two dimensional nature of the Fermi surface. However, for the critical boson $\varphi$, its quasi two-dimensional dispersion leads to important consequences. To be more concrete, the computation of the energy flux later relies on the matrix elements summed over polarizations, which we can be simplified based on Goldstone’s theorem $\{\sum_{i}|N^{i}_{\bm{q}}|,~{}\sum_{i}|L^{i}_{\bm{k},\bm{q}}|\}\sim A_{N,L}({\left|\bm{q_{\text{2D}}}\right|}^{2}+\lambda q_{z}^{2})$. Here, $A_{N}$ and $A_{L}$ are proportionality constants with the bare dimension of $[\text{energy}]^{2}$ and $[\text{energy}]^{0}$, respectively. The dependence on phonon momentum $\bm{q}$ can be decomposed into its 2D projection $\bm{q}_{\text{2D}}$ and the $z$-component $q_{z}$. The ${\left|\bm{q_{\text{2D}}}\right|}^{2}$ part arises from in-plane strain and the $q_{z}^{2}$ part arises from out-of-plane strain. We will assume that $\lambda\ll 1$ so that the $q_{z}^{2}$ term is a small contribution. We will comment later on the effects of larger $\lambda$.
As emphasized previously, we will assume that ${\mathcal{L}}_{ep}$ is weak compared to the self-interactions within the electron and the phonon sectors, allowing for a perturbative computation of the energy flux $\kappa$ between the two sectors. It is worth noting that the linear hybridization in ${\mathcal{L}}_{ep}^{(2)}$ can nominally be eliminated by an appropriate redefinition and change of basis of the $\varphi$-boson. However, we implicitly assume here that the phonons are responsible for heat dissipation to the environment. Eliminating Eq. (4b) via a basis transformation will then introduce a coupling between $\varphi$ and the environment, which we would like to avoid. Therefore, we explicitly include the weak linear hybridization in Eq. (4b) and keep track of how heat is dissipated away from the purely electronic components to the phonons, and subsequently to the environment.
A summary of our results is presented in Fig. 2, which are obtained using the Keldysh formalism [79]. The different contributions, $\kappa^{(\ell)}$, correspond to the energy flux due solely to the couplings, ${\mathcal{L}}_{ep}^{(\ell)}$ ($\ell=1,2,3$). All of these individual contributions add up together to yield the total $\kappa$, and $\Gamma_{E}$ is then determined via Eq. (2). In panel (a), we show the contribution due to the conventional electron-phonon coupling Eq. (4a), which is similar to Allen’s classic result [31] with a single crossover temperature scale around the Debye frequency ${\omega_{\rm D}}$. In panel (b), we show the additional structure in the $T$-dependent crossovers due to the boson-phonon couplings Eq. (4b) and (4c), which will be the subject of our discussion below.
Figure 2: (a) Numerical plot of electronic contribution $\kappa^{(1)}$ to the energy flux. (b)Different regimes of the energy flux due to phonon-boson coupling (Eqs. (4b),(4c)) near the quantum critical point. $\omega_{D}$ is the Debye frequency of the phonons; $c$ is the speed of sound; $v_{\varphi}$ is the velocity of the $\varphi$ boson; $\gamma$ is the Landau damping coefficient (see Eq. (7)); $m_{0}$ is the renormalized boson mass which vanishes at the quantum critical point.
Energy-relaxation due to deformation potential coupling.- To make analytical progress in evaluating the energy flux, we have made some simplifying physical approximations, where we assume that (i) energy relaxation occurs homogeneously in space, (ii) the phonons remain reasonably well-defined quasiparticles, a point we revisit later, and in quasi-equilibrium at a temperature, $T_{p}$ [79]. The leading contribution due to ${\mathcal{L}}_{ep}^{(1)}$ in Eq. (4a) leads to the phonon self-energy in Fig. 1(b). When $T_{e}\approx T_{p}\approx T$, we obtain $\partial_{t}E=\kappa^{(1)}(T_{e}-T_{p})$, where the general expression for $\kappa^{(1)}$ is quite involved [79]. However, neglecting the electron self-energies and assuming $c\ll v_{F}$, we obtain
$\kappa_{\text{free}}^{(1)}=\frac{{\mathcal{N}}}{4v_{F}a_{z}}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\overline{|M({\left|\bm{q}\right|})|^{2}}\frac{c^{2}{\left|\bm{q}\right|}^{2}}{{\left|\bm{q_{\text{2D}}}\right|}}\frac{1}{T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\,,$ (5)
where $\bm{q}$ is the momentum transfer between the phonons and the particle-hole excitations, $a_{z}$ is the lattice constant in the $z-$direction and ${\mathcal{N}}$ is the electronic density of states. The $1/{\left|\bm{q_{\text{2D}}}\right|}$ factor arises from projecting onto the 2D Fermi surface, which however does not lead to any IR divergence. We have also replaced $\sum_{i}|M^{i}_{\bm{k}+\bm{q/2},\bm{q}}|^{2}$ by its average over the Fermi surface $\overline{|M(\bm{q})|^{2}}$. When $T\ll{\omega_{\rm D}}$ (Debye frequency), each factor of ${\left|\bm{q}\right|}$ and ${\left|\bm{q_{\text{2D}}}\right|}$ is replaced by $T$ (recall that $|M({\left|\bm{q}\right|})|^{2}\propto{\left|\bm{q}\right|}^{2}$), and we obtain $\kappa^{(1)}_{\text{free}}\propto T^{4}$. This result can also be alternatively derived using Fermi’s Golden rule, as follows:
$\begin{split}&\kappa^{(1)}\sim\overbrace{{\mathcal{N}}}^{\text{Fermion DOS}}\times\overbrace{T^{3}}^{\text{phonon density}}\times\overbrace{(\sqrt{T})^{2}}^{\text{Goldstone coupling}}\\ &\times\underbrace{T}_{\text{Energy transfer per particle}}\times\underbrace{\frac{1}{T}}_{\text{Expand around equilibrium}}\,.\end{split}$ (6)
Here a typical phonon has energy $T$ and the coupling enters as $\sqrt{T}$ because the displacement field $X$ is related to the creation and annihilation operators via $X_{\bm{q}}=(1/\sqrt{2c{\left|\bm{q}\right|}})(a_{\bm{q}}+a_{-\bm{q}}^{\dagger})$ [80]. On the other hand, when $T\gg{\omega_{\rm D}}$, Eq. (5) becomes $T$-independent, which is the classic expectation that in the equipartition regime of the phonons, it becomes difficult for them to absorb energy from the electrons. The numerical result of Eq. (5) is plotted in Fig. 2(a), which reproduces Allen’s result [31]; upon including the marginal Fermi liquid self-energy, the results remain nearly identical to the Fermi liquid results except at very low-temperatures [79].
Energy-relaxation due to linear coupling between phonon and collective mode.- Let us now turn to the contribution due to ${\mathcal{L}}_{ep}^{(2)}$ in Eq. (4b), with a phonon self-energy shown in Fig. 1(c). As noted previously, the renormalized propagator for $\varphi$ in the MFL is overdamped with $z_{\varphi}=2$, with a spectral function,
$A_{\varphi}(\omega,\bm{q})=\frac{-2\gamma\omega}{\gamma^{2}\omega^{2}+[{\left|\bm{q_{\text{2D}}}\right|}^{2}v_{\varphi}^{2}+m^{2}(T)]^{2}}\,.$ (7)
Here, $\gamma={\mathcal{N}}g^{2}/\Gamma$ is the Landau-damping coefficient with dimension of energy (${\mathcal{N}}$ is the fermion density of states, $\Gamma$ is the fermion elastic scattering rate), which we expect to be an electronic energy scale. The boson mass term, $m^{2}(T)$ also includes the thermal corrections due to a finite temperature, that are potentially important in the quantum critical regime. The QCP corresponds to $m^{2}(T=0)=0$, and the thermal fluctuations induce a $m^{2}(T)\sim\gamma T\ln(\gamma/T)$ [71, 81].
Proceeding as before, and assuming $\varphi$ well equilibrated with the electrons at $T_{e}$, with $T_{e}\approx T_{p}\approx T$, we obtain $\partial_{t}E=\kappa^{(2)}(T_{e}-T_{p})$, where
$\kappa^{(2)}=\frac{1}{4}\sum_{i}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\left|N^{i}_{\bm{q}}\right|^{2}\left(-A_{\varphi}(c{\left|\bm{q}\right|},\bm{q})\right)\frac{c{\left|\bm{q}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\,.$ (8)
Compared to Eq. (5), the contribution from the particle-hole bubble is replaced by the $\varphi$-boson spectral function. To analyze the integral for $T\ll{\omega_{\rm D}}$, we make the following simplifications: (1) The typical phonon momentum ${\left|\bm{q}\right|}$ is replaced by $T/c$. (2) We approximate the matrix element by Goldstone’s theorem $\sum_{i}|N_{\bm{q}}^{i}|^{2}\to A_{N}{\left|\bm{q_{\text{2D}}}\right|}^{2}$, and we return to the neglected $q_{z}$ component later. With these, the integral simplifies to
$\kappa^{(2)}\approx\frac{A_{N}}{c^{5}}\int_{0}^{\pi}{\rm d}\theta\frac{\gamma T^{5}\sin^{3}\theta}{\gamma^{2}T^{2}+\left[m^{2}(T)+(\sin^{2}\theta)T^{2}v_{\varphi}^{2}/c^{2}\right]^{2}}\,.$ (9)
Here $\theta$ is the angle between phonon momentum $\bm{q}$ and the $z$-axis [79].
$\kappa^{(2)}$ exhibits a complex sequence of crossovers marked by different power-laws as a function of temperature, as illustrated in Fig. 2 (b) and Fig. 3(a). They are as follows:
(i) Let us start slightly away from the QCP, $m(T=0)=m_{0}>0$, which is assumed to be smaller than all other energy scales, including ${\omega_{\rm D}}$. At the lowest temperature, the denominator of Eq. (9) is dominated by $m_{0}$, so we obtain $\kappa^{(2)}\sim(A_{N}/c^{5})(\gamma/m_{0}^{4})T^{5}$. In this regime, $\varphi$ mediates local interaction between the electrons, and the effects of damping and their dispersion are suppressed relative to $m_{0}$. Alternatively, by counting the phonon phase-space, energy transfer and expansion around the thermal equilibrium, we can obtain $\kappa^{(2)}\sim A_{N}T^{4}\times\braket{\varphi\varphi}$. In this low-energy and long-wavelength limit, the boson $\varphi$ is expected to behave roughly as an ohmic bath, which yields $\braket{\varphi\varphi}\propto\gamma T/m_{0}^{4}$. We also remark in passing that these bosonic collective modes are generically present in interacting Fermi liquids, and $\kappa^{(2)}$ can be interpreted as an interaction correction to energy relaxation on top of the more dominant free fermion result $\kappa^{(1)}$. When the system is deep in the Fermi liquid phase, the boson mass $m_{0}$ is expected to be an electronic energy scale that overwhelms other crossovers, i.e. $\kappa^{(2)}$ grows as $T^{5}$ until being cutoff at $T\sim{\omega_{\rm D}}$.
(ii) With increasing $T>m_{0}^{2}/\gamma$, the $\gamma^{2}T^{2}$ term in the denominator of Eq. (9) dominates, so $\kappa^{(2)}\sim(A_{N}/c^{5})T^{3}/\gamma$, which becomes comparable to $\kappa^{(1)}$, albeit with a different coefficient. In this regime, the $z_{\varphi}=2$ dynamics first rears its head; however, the typical phonon momentum is small compared to the typical $\varphi$ frequencies. Therefore, the phonons are more sensitive to the homogeneous fluctuations of $\varphi$, as opposed to their full diffusive character. Mathematically, regimes (i) and (ii) can be described in a unified fashion by defining $\Delta(T)=\max(m(T),\sqrt{\gamma T})$ and writing $\kappa^{(2)}\sim(A_{N}/c^{5})\gamma T(T/\Delta(T))^{4}$; this applies to the quantum critical fan ($m_{0}=0$) as well as for $\Delta(T)\to m(T)$.
(iii) When $T\gtrsim\gamma(c^{2}/v_{\varphi}^{2})$, the $T^{2}v_{\varphi}^{2}/c^{2}$ term in the denominator becomes relevant. The $\theta$-integral in Eq. (9) should be cutoff at $\theta_{\text{min}}\sim(c/v_{\varphi})\Delta(T)/T$, which yields $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma T\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$, which signals an enhanced contribution due to phonons moving in the $z$ direction. In this regime, the typical phonon momentum is large compared to the typical $\varphi$ frequencies, and unlike the previous lower-temperature regime, the phonons are now sensitive to diffusive character of $\varphi$. This regime is unique to phonons for the following reason. Suppose if we were to replace the phonons by a distinct bosonic mode of electronic origin, but now with $c\approx v_{\varphi}$. Then this regime would onset at $T\sim\gamma c^{2}/v_{\varphi}^{2}\approx\gamma$, which is a high electronic energy scale. In the present setting associated with phonons, the onset of the $\kappa^{(2)}\sim T\ln(T)$ regime is tied to the parametric suppression in $c\ll v_{\varphi}$.
(iv) With a further increase $T\gtrsim{\omega_{\rm D}}$, the integral saturates, and we obtain a constant $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma\omega_{D}\ln(v_{\varphi}^{2}\omega_{D}/(c^{2}\gamma))$.
It is useful to make a few remarks about the underlying dimensionality of the problem. The lowest temperature regimes (i), (ii) exhibiting $\kappa^{(2)}\sim T^{5}$ and $\kappa^{(2)}\sim T^{3}$ depend on the phonon dimensionality in a straightforward fashion via their density of states. However, the $\kappa^{(2)}\sim T\ln(T)$ regime (iii) involves a subtle kinematic interplay between the phonon and $\varphi$. In particular, the phonon dispersion along the $z-$direction contributes more significantly compared to the other directions, leading to the $\ln$ enhancement; the latter would be absent if the phonons were purely two-dimensional. Furthermore, $\kappa^{(2)}$ in this regime also depends crucially on the assumption of completely decoupled nature of the two-dimensional electronic layers, via the simplification, $N^{i}_{\bm{q}}\propto{\left|\bm{q_{\text{2D}}}\right|}$, instead ${\left|\bm{q}\right|}$. Finally, let us also consider the additional perturbation due to coupling the $\varphi$ boson to out-of-plane strain, which leads to a correction, $\lambda q_{z}^{2}$ term in $|N^{i}_{\bm{q}}|^{2}$ with $\lambda\ll 1$. This does not lead to qualitatively different results in regimes (i) and (ii), but it can enhance the $\ln$ factor in regime (iii) to a linear factor of $v_{\varphi}^{2}T/(c^{2}\gamma)$ [79]. Since this regime is upper bounded by $T\sim\omega_{D}$, this term remains a small perturbation if $\lambda\ll\gamma c^{2}/(\omega_{D}v_{\varphi}^{2})$.
Figure 3: Numerical plot of (a) $\kappa^{(2)}(T)$ and (b) $\kappa^{(3)}(T)$ on log-log scale. The blue curves are numerical evaluation of Eqs. (8) and (10). The dashed lines denote the approximate scaling forms deduced analytically.
Energy-relaxation due to non-linear coupling between phonon and collective mode.- Finally, we consider the contribution due to ${\mathcal{L}}_{ep}^{(3)}$ in Eq. (4c), with a phonon self-energy shown in Fig. 1(d), which involves scattering between $\varphi^{2}$ and the acoustic phonons. The renormalized spectral function for $\varphi$ is still given by Eq. (7). The general expression for the flux $\kappa^{(3)}$ reads
$\begin{split}&\kappa^{(3)}=-\frac{1}{16}\int_{\bm{q}}\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\frac{c{\left|\bm{q}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\int\frac{{\rm d}\nu}{2\pi}\int\frac{{\rm d}^{3}\bm{k}}{(2\pi)^{3}}\\ &\times\sum_{i}|L^{i}_{\bm{k}+\bm{q}/2,\bm{q}}|^{2}A_{\varphi}(\nu,\bm{k})A_{\varphi}(\nu+c{\left|\bm{q}\right|},\bm{k}+\bm{q})\\ &\times\left[\coth\bigg{(}\frac{\nu+c{\left|\bm{q}\right|}}{2T}\bigg{)}-\coth\bigg{(}\frac{\nu}{2T}\bigg{)}\right]\,.\end{split}$ (10)
Once again, there is a complex sequence of temperature dependent crossovers that can be obtained on the basis of following approximations [79]: replace $\sum_{i}|L^{i}_{\bm{k}+\bm{q}/2,\bm{q}}|^{2}$ by $A_{L}{\left|\bm{q_{\text{2D}}}\right|}^{2}$, and for $T\lesssim{\omega_{\rm D}}$, include contributions from phase-space where $c{\left|\bm{q}\right|}\sim\nu\sim T$. Then $\kappa^{(3)}$ is schematically given by,
$\begin{split}\kappa^{(3)}&\approx\frac{A_{L}}{a_{z}c^{5}}T^{5}\int_{0}^{\pi}{\rm d}\theta\sin^{3}\theta\int{\rm d}^{2}\bm{k}_{\text{2D}}\\ &\times A_{\varphi}(T,\bm{k}_{\text{2D}})A_{\varphi}(T,\bm{k}_{\text{2D}}+\bm{q}_{\text{2D}}(\theta))\,,\end{split}$ (11)
where the vector $\bm{q}_{\text{2D}}(\theta)$ has norm $(T/c)\sin\theta$, and $\theta$ is the angle between $\bm{q}$ and the z-axis.
The regimes where $\kappa^{(3)}$ exhibits the distinct $T-$dependent crossovers are identical to the ones that arise in the discussion for $\kappa^{(2)}$ above. (i) For $T\lesssim m_{0}^{2}/\gamma$, the $\varphi$-spectral functions are dominated by $m_{0}$ and the $\bm{k}$-integral can be estimated as $A_{\varphi}^{2}(T,k_{\text{typ}})k_{\text{typ}}^{2}$ treating the $\bm{q}_{\text{2D}}(\theta)$ term as being unimportant. Here $k_{\text{typ}}$ is the typical magnitude of $\bm{k}$, estimated to be $k_{\text{typ}}\sim m_{0}/v_{\varphi}$. Therefore, we obtain $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})\gamma^{2}T^{7}/m_{0}^{6}$, which is strongly suppressed compared to $\kappa^{(2)}$, as one might expect based on scaling arguments. Alternatively, a simple counting argument leads to $\kappa^{(3)}\sim A_{L}T^{4}\times\braket{\varphi^{4}}$. Here $\braket{\varphi^{4}}\propto T^{3}/m_{0}^{6}$ where the $T^{2}/m_{0}^{8}$ comes from two ohmic baths and there are $N\propto Tm_{0}^{2}$ different ways to distribute the energy between and the momentum these two baths. (ii) As before, when $m_{0}^{2}/\gamma<T<\gamma(c^{2}/v_{\varphi}^{2})$, we can replace $m_{0}^{2}$ by $\gamma T$, and obtain $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})T^{4}/\gamma$. If we take $m_{0}\rightarrow 0$ in the quantum critical region, these regimes crossover smoothly into each other with $\kappa^{(3)}\sim(A_{L}/a_{z}c^{5}v_{\varphi}^{2})\gamma^{2}T^{7}/m^{6}(T)$.
(iii) When $\gamma(c^{2}/v_{\varphi}^{2})<T<{\omega_{\rm D}}$, the $\bm{q}_{\text{2D}}(\theta)$ term in the second $A_{\varphi}(...)$ factor becomes more important than $\bm{k}_{\text{2D}}$, leading to $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})\gamma T^{2}\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$. The logarithmic factor has the same origin as that in $\kappa^{(2)}$. In the quantum critical regime, this becomes $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})(\gamma^{2}T^{3}/m^{2}(T))\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$. Finally, when $T>{\omega_{\rm D}}$, $\kappa^{(3)}\sim(A_{L}/a_{z}cv_{\varphi}^{6})\gamma T{\omega_{\rm D}}\ln(v_{\varphi}^{2}{\omega_{\rm D}}/(c^{2}\gamma))$. Here, the high-$T$ limit of $\kappa^{(3)}$ is different from $\kappa^{(1)}$ and $\kappa^{(2)}$ in that it does not saturate but grows linearly in $T$. This is expected as $\kappa^{(3)}$ involves an additional boson compared to $\kappa^{(2)}$, and therefore the result should be proportional to the density of that additional boson, which is proportional to $T$, and this will be ultimately cut off at an electronic energy scale where $\varphi$ reveals its electronic nature. A comparison between analytical estimate and numerical result is shown in Fig. 3(b).
Connection to Experiments.- Recent optical experiments involving the technique of two-dimensional coherent spectroscopy [82, 83, 84] have studied the nonlinear response of correlated metals and its relationship to energy relaxation [85, 39], providing a further impetus to a careful theoretical examination of the questions raised above. Over a broad range of hole-dopings in a specific cuprate family [86, 87], recent pump-probe measurements [39] have identified energy relaxation as the dominant nonlinear process that governs the experimental phenomenology. At low temperatures across most of the doping phase diagram, $\Gamma_{E}\sim k_{B}T/\hbar$, with a rate that is significantly slower (approximately by an order of magnitude) compared to the Planckian momentum-relaxation rate [1, 88]. With increasing temperature, there is a tendency for $\Gamma_{E}$ to saturate across a doping dependent crossover scale. Both of these regimes are in stark contrast to the results expected in a conventional Fermi liquid coupled to acoustic phonons [31]. In THz experiments, it is the second rate $\Gamma_{ep}$ that is measured, which leads to a fast initial decay of the pump-probe signal to a plateau at the order of pico seconds, whereas the phonon-environment relaxation $\Gamma_{p-env}$ leads to a slower decay on longer timescales.
To make contact with experiments, we can convert $\kappa=[\kappa^{(1)}+\kappa^{(2)}+\kappa^{(3)}]$ to the energy relaxation rate, $\Gamma_{E}$, via Eq. (2). Here, due to the comparison between the heat capacities $C_{p}\sim(T/{\omega_{\rm D}})^{3}$ and $C_{e}\sim(T/E_{F})\ln(\gamma/T)$, an additional crossover is present, whose temperature scale $T_{*}\sim\sqrt{{\omega_{\rm D}}^{2}/E_{F}}$ is set by equating $C_{p}$ and $C_{e}$. Eq. (2) states that the sector with the smaller heat capacity dominates as its temperature changes more rapidly. When $T\ll T_{*}$, we have $\Gamma_{E}\approx\kappa/C_{p}$ and when $T\gg T_{*}$, $\Gamma_{E}\approx\kappa/C_{e}$. From thermodynamic measurements in a representative family of the cuprates [89], $T_{*}$ is at the order of $1$-$10$ K. For the experimental range of $T\sim O(100\textrm{ K})$, the latter is more relevant, so we would then divide the scaling of $\kappa$ by a power of $T$ to get scaling of $\Gamma_{E}$. Our theoretical results for $\kappa^{(3)}$ provide an appealing scenario for the experimental observations. In the experimental temperature range, the energy relaxation rate $\Gamma_{E}$ is related to $\kappa^{(3)}$ by $\Gamma_{E}\approx\kappa^{(3)}/C_{e}$. Therefore, in regime (iii) and regime (iv) we would obtain a crossover from $T\ln T$ to $T$-independent constant in $\Gamma_{E}$, which qualitatively agrees with the observation in [39] for samples near optimal doping (see Fig. 6c in [39]).
We also comment briefly on the momentum-drag phenomenon [90, 91, 17, 19], which plays a potentially important role in electrical transport in clean systems. In a translational invariant system, the momentum injected into the system is passed around rapidly between the electronic and (critical) bosonic subsystems via rapid collisions, while the system as a whole conserves total momentum. Electrical transport is then determined by the dissipation rate of the total momentum. However, the energy relaxation problem we study in this work is different, since we assume that the phenomenology is dominated by how fast the energy is passed from the hot electron sector to the cold phonon sector, and the dissipation of the total energy to the environment happens at a longer time scale. Therefore, if we model the electron and the phonon as an isolated system, the physics of interest happens before the time scale of drag effects.
Outlook.- We have demonstrated here that in the vicinity of an electronic quantum critical point, the symmetry-allowed coupling between the electronic degrees of freedom and acoustic phonons can lead to a set of complex temperature-dependent crossovers in the energy relaxation rate. When the momentum relaxation in a marginal Fermi liquid is controlled primarily by electronic mechanisms and its interplay with disorder, while the energy relaxation is limited by the flux to the phononic bath, there is no a priori reason for the two relaxation rates to exhibit a similar behavior. Indeed, our theoretical results have unambiguously demonstrated that energy relaxation proceeds via a multitude of crossovers and temperature dependencies across emergent low-energy scales that are distinct from any momentum relaxation rate.
In recent years, inspired in part by experiments [92, 93, 94], there have been suggestions that even the phonons in a strongly correlated metal can lose their quasiparticle-like character [90, 95, 96, 97, 91]. The common underlying theme is that the phonon lifetime is also Planckian, $O(\hbar/k_{B}T)$. Our theoretical conclusions for the energy flux should broadly remain applicable to these incoherent phonons as well in spite of broadened phonon “quasiparticle-peak”, since their typical frequency is still of order $k_{B}T$. A detailed theoretical study of an entangled liquid made out of a non-Fermi liquid of electrons and phonons is an interesting problem for the future.
In this work, while we have focused only on the normal state and ignored the low-temperature superconducting phase, it is likely that the pairing fluctuations above superconducting $T_{c}$ as well as the correlated Bogoliubov quasiparticles below $T_{c}$ , will affect the energy relaxation processes via their coupling to phonons [98], which remains an exciting topic for future research. Finally, given the complex set of crossovers in the energy-relaxation demonstrated by our results near the quantum critical point due to the mismatch between the velocities of the phonons and collective modes, and their dimensionality, we anticipate similar regimes to also appear in the sound attenuation [99, 100, 101, 102, 103, 104, 34]. We leave a detailed analysis of sound attenuation in quantum critical metals for future investigation.
Acknowledgments.- We thank N.P. Armitage, D. Chaudhuri, and B. Ramshaw for stimulating discussions. H.G. is supported by a Wilkins postdoctoral fellowship at Cornell University. D. C. is supported in part by a NSF CAREER grant (DMR-2237522), and a Sloan Research Fellowship.
Supplementary Material for “Phonon Induced Energy Relaxation in Quantum Critical Metals”
Haoyu Guo and Debanjan Chowdhury
I Keldysh formalism for energy relaxation
In this section, we will set up a Keldysh field-theoretic formalism [105, 106] to compute the energy transfer rate $\kappa$ between the electron and the phonon sectors, respectively. Within the scope of the two-temperature model [31], we ignore the dissipation to the environment such that the total energy of the electrons and the phonons is conserved. For simplicity, we compute the energy gain of the phonon sector. Using the phonon-only Lagrangian (Eq.(3b)), we apply Noether’s theorem and obtain the Hamiltonian of the phonon sector,
${\mathcal{H}}=\int\sum_{\bm{q}}{\rm d}t\frac{1}{2}\left[\partial_{t}X_{i,-\bm{q}}\partial_{t}X_{i,\bm{q}}+c^{2}\bm{q}^{2}X_{i,-\bm{q}}X_{i,\bm{q}}\right]\,.$ (12)
Note that we have Wick rotated from the imaginary-time formulation to real-time formulation using $t=-i\tau$. Computing the expectation value of ${\mathcal{H}}$ amounts to replacing $X_{i}X_{i}$ to the greater and lesser Green’s function $D^{\gtrless}_{X,ij}$ of the $X_{i}$ field, which is formally defined as
$D^{>}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $i\braket{X_{i}(t,\bm{x})X_{j}(t^{\prime},\bm{x}^{\prime})}\,,$ (13a)
$D^{<}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $i\braket{X_{j}(t^{\prime},\bm{x}^{\prime})X_{i}(t,\bm{x})}\,.$ (13b)
We further define the retarded, advanced and the Keldysh Green’s function using
$D_{R,ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $[D^{>}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})-D^{<}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})]\theta(t-t^{\prime})\,,$ (14a)
$D_{A,ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $[D^{>}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})-D^{<}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})](-\theta(t^{\prime}-t))\,,$ (14b)
$D_{K,ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $D^{>}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})+D^{<}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})\,,$ (14c)
where $\theta$ is the step function. Next, the spectral function is defined as
$A_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})=i[D_{R,ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})-D_{A,ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})]=i[D^{>}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})-D^{<}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})]\,.$ (15)
We parameterized the Keldysh Green’s function using the distribution function $F_{X}$:
$D_{K}=D_{R}\circ F_{X}-F_{X}\circ D_{A}\,,$ (16)
where $\circ$ means convolution in both space-time and cartesian indices. Finally, we consider the Wigner transform of the above two-point functions, which is the fourier transform in the relative coordinate ($F$ below is an arbitrary two-point function)
$F(\omega,\bm{k};t,\bm{x})=\int{\rm d}\delta t{\rm d}^{3}\delta\bm{x}\exp(-i\bm{k}\cdot\delta\bm{x}+i\omega\delta t)F(t+\delta t/2,\bm{x}+\delta\bm{x}/2;t-\delta t/2,\bm{x}-\delta\bm{x}/2)\,.$ (17)
In a quasi-equilibrium state with temperature $1/\beta$, the Wigner transform of the distribution function is $\coth(\beta\omega/2)$ for boson and $\tanh(\beta\omega/2)$ for fermion.
With the definitions above, the energy density $E=\braket{{\mathcal{H}}}$ can be written in terms of the Wigner transform of spectral function and the distribution function as
$\begin{split}E(t,\bm{x})=-\sum_{\bm{q}}\int\frac{{\rm d}\Omega}{2\pi}\frac{\Omega^{2}+c^{2}\bm{q}^{2}}{2}\times\sum_{i,j}\frac{F_{X,ij}(\Omega,\bm{q};t,\bm{x})-\delta_{ij}}{2}A_{ji}(\Omega,\bm{q};\bm{x},t)+\dots\,.\end{split}$ (18)
Here $\dots$ represent the higher gradient terms arising from the Wigner transform.
We will compute $\partial_{t}E(t,\bm{x})$ by making the following simplifications: (1) We assume that energy relaxation is homogeneous in space and slow, so we drop the explicit dependence on $\bm{x}$, together with higher gradient terms ($\dots$). (2) We assume the phonons remain well-defined quasiparticles with the following spectral function :
$A_{ij}(\Omega,\bm{q})=-\frac{2\pi{\rm sgn\,}\Omega}{2c{\left|\bm{q}\right|}}\left[\delta(\Omega+c{\left|\bm{q}\right|})+\delta(\Omega-c{\left|\bm{q}\right|})\right]\delta_{ij}\,.$ (19)
(3) The phonon sector is in quasi-equilibrium, described by the distribution function $F_{X,ij}(\Omega,\bm{q};t)=\delta_{ij}F_{X}(\Omega;t)$ where $F_{X}(\Omega;t)=\coth(\Omega/2T_{p})$ with $T_{p}$ the phonon temperature that varies with time slowly. (4) Finally, the time-dependence mainly arises from $\partial_{t}F$, which is described by the (spatially homogeneous) Keldysh kinetic equation [106]
$\begin{split}2\Omega\partial_{t}F_{X,ij}(\Omega,\bm{q};t)=iF_{X}(\Omega;t)\left[\Pi_{R,ij}(\Omega,\bm{q})-\Pi_{A,ij}(\Omega,\bm{q})\right]-i\Pi_{K,ij}(\Omega,\bm{q})\,,\end{split}$ (20)
where $\Pi_{R/A/K}$ denotes the retarded, advanced, and the Keldysh self-energy of the phonon, all evaluated in the quasi-equilibrium state. Assembling everything, we obtain the following expression for the energy transfer rate
$\begin{split}\partial_{t}E=\frac{i}{4}\sum_{\bm{q},i}\left\{F_{X}(\Omega;t)\left[\Pi_{R,ii}(c{\left|\bm{q}\right|},{\left|\bm{q}\right|})-\Pi_{A,ii}(c{\left|\bm{q}\right|},{\left|\bm{q}\right|})\right]-\Pi_{K,ii}(c{\left|\bm{q}\right|},{\left|\bm{q}\right|})\right\}\,.\end{split}$ (21)
This is the main result of this subsection. Strictly speaking, Eq.(21) does not include all contributions to the energy as we have the ignored the contribution due to the electron-phonon couplings. However, these effects are assumed to be higher order perturbative corrections in terms of the weak electron-phonon couplings, so we ignore them for now.
The spectral function (19) sets the phonon frequency $\Omega$ to be exactly equal to dispersion $c{\left|\bm{q}\right|}\approx T$ (temperature). When the phonons become incoherent, $A_{ij}(\Omega,\bm{q})$ is replaced by a broader Lorentzian function with width of order $T$, and as a result the typical frequency $\Omega$ is still at the same order, which from the perspective of power-counting estimate serves the same purpose.
Next we evaluate Eq.(21) for the three couplings introduced in the main text. We will begin by first considering the effect of ${\mathcal{L}}_{ep}^{(1)}$ in Eq. (4a) on energy relaxation, thus reproducing and generalizing the previously known results [31], before presenting the new results due to ${\mathcal{L}}_{ep}^{(2)}+{\mathcal{L}}_{ep}^{(3)}$ in Eq. (4b)-(4c).
I.1 Deformation potential coupling
We start by evaluating the energy transfer rate due to the direct electron-phonon deformation potential coupling, ${\mathcal{L}}_{ep}^{(1)}$, in Eq. (4a). The associated phonon self-energy is given by
$\Pi^{\gtrless}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})=i\hat{M^{i}}(\bm{x},\bm{x}^{\prime})\hat{M^{j}}(\bm{x},\bm{x}^{\prime})G^{\gtrless}(t,\bm{x};t^{\prime},\bm{x}^{\prime})G^{\lessgtr}(t^{\prime},\bm{x}^{\prime};t,\bm{x})\,.$ (22)
Here $G$ is the fermion Green’s function and $\hat{M}^{i}$ is the fourier transform of the matrix element $M^{i}_{\bm{k},\bm{q}}$ to real space, which only serves for a book-keeping purpose and will be converted back to momentum space below. The explicit definition of $G$ is
$G^{>}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $-i\braket{c(t,\bm{x})c^{\dagger}(t^{\prime},\bm{x}^{\prime})}\,,$ (23a)
$G^{<}(t,\bm{x};t^{\prime},\bm{x}^{\prime})$ $=$ $i\braket{c^{\dagger}(t^{\prime},\bm{x}^{\prime})c(t,\bm{x})}\,.$ (23b)
The other Green’s functions are derived from these similar to Eqs. (14)-(16).
We convert the self-energies to other components using Eq. (14), and also transform them to momentum space. Since we assume a quasi-equilibrium state, the Keldysh Green’s function is related to the spectral function by
$G_{K}(\omega,\bm{k})=-iA(\omega,\bm{k})F_{c}(\omega)\,,$ (24)
and similar relations hold for the phonons. The energy transfer rate from Eq. (21) reads
$\partial_{t}E=\frac{1}{4}\sum_{\bm{q},\bm{k}}\int\frac{{\rm d}\nu}{2\pi}\sum_{i}|M^{i}_{\bm{k+q/2},\bm{q}}|^{2}A_{c}(\nu,\bm{k})A_{c}(\nu+c{\left|\bm{q}\right|},\bm{k}+\bm{q})\left[F_{X}(c{\left|\bm{q}\right|})\left(F_{c}(c{\left|\bm{q}\right|}+\nu)-F_{c}(\nu)\right)+F_{c}(c{\left|\bm{q}\right|}+\nu)F_{c}(\nu)-1\right]\,,$ (25)
where, $F_{c}(\omega)$ is the fermion distribution at $T=T_{e}$ (local equilibrium of the electron sector), and $A_{c}(\omega,\bm{k})$ is the fermion spectral function,
$F_{c}(\omega)$ $=$ $\tanh\bigg{(}\frac{\omega}{2T_{e}}\bigg{)}\,,$ (26a)
$A_{c}(\omega,\bm{k})$ $=$ $\frac{-2\Sigma^{\prime\prime}(\omega)}{\left(\omega-\Sigma^{\prime}(\omega)-\xi_{\bm{k}}\right)^{2}+\left(\Sigma^{\prime\prime}(\omega)\right)^{2}}\,.$ (26b)
Here $\xi_{\bm{k}}=(\varepsilon_{\bm{k}}-\mu)$ is the electronic dispersion measured from the Fermi surface, and we have assumed the (retarded) fermion self-energy $\Sigma(\omega)=\Sigma^{\prime}(\omega)+i\Sigma^{\prime\prime}(\omega)$ to be approximately independent of $\xi_{\bm{k}}$ near the quantum critical point [71, 19]. We will employ the marginal Fermi liquid ansatz [77] by setting the imaginary part $\Sigma^{\prime\prime}(\omega)=-\Gamma/2-\alpha|\omega|$, where $\Gamma$ is the elastic scattering rate due to disorders, and $\alpha$ is a dimensionless coefficient related to fermion-boson Yukawa coupling and disorder strength [19, 18]. The real part $\Sigma^{\prime}(\omega)$ is related via the Kramers-Krönig relation. Linearizing Eq. (25) for $T_{e}\approx T_{p}\approx T$, we obtain $\partial_{t}E=\kappa^{(1)}(T_{e}-T_{p})$, where
$\kappa^{(1)}=\frac{1}{8}\sum_{i}\int\frac{{\rm d}\nu}{2\pi}\int\frac{{\rm d}^{3}\bm{q}{\rm d}^{3}\bm{k}}{(2\pi)^{6}}|M^{i}_{\bm{k+q/2},\bm{q}}|^{2}A_{c}(\nu,\bm{k})A_{c}(\nu+c{\left|\bm{q}\right|},\bm{k}+\bm{q})\frac{c{\left|\bm{q}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\left[\tanh\bigg{(}\frac{\nu+c{\left|\bm{q}\right|}}{2T}\bigg{)}-\tanh\bigg{(}\frac{\nu}{2T}\bigg{)}\right]\,.$ (27)
To make the kinematic contributions to Eq. (27) transparent, we make the following further simplifications. We assume the Fermi surface is isotropic, and replace the interaction matrix element $\sum_{i}|M^{i}_{\bm{k}+\bm{q/2},\bm{q}}|^{2}$ by its average over the Fermi surface $\overline{|M(\bm{q})|^{2}}$, enabling an analytical computation that leads to,
$\begin{split}&\kappa^{(1)}=\frac{1}{8a_{z}}\int\frac{{\rm d}^{3}\bm{q}}{(2\pi)^{3}}\int\frac{{\rm d}\nu}{2\pi}{\mathcal{N}}^{2}\frac{4\pi^{2}}{k_{F}}\overline{|M(\bm{q})|^{2}}\times\left[\frac{1}{\sqrt{{\left|\bm{q_{\text{2D}}}\right|}^{2}+(A+iB)^{2}}}+\frac{1}{\sqrt{{\left|\bm{q_{\text{2D}}}\right|}^{2}+(A-iB)^{2}}}\right]\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\frac{c{\left|\bm{q}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|\bm{q}\right|}}{2T}}\left[\tanh\bigg{(}\frac{\nu+c{\left|\bm{q}\right|}}{2T}\bigg{)}-\tanh\bigg{(}\frac{\nu}{2T}\bigg{)}\right]\,,\end{split}$ (28)
where $A,B$ are given by
$A$ $=$ $-\frac{\Sigma^{\prime\prime}(\nu)+\Sigma^{\prime\prime}(\nu+c{\left|\bm{q}\right|})}{v_{F}}\,$ (29a)
$B$ $=$ $\frac{c{\left|\bm{q}\right|}+\Sigma^{\prime}(\nu)-\Sigma^{\prime}(\nu+c{\left|\bm{q}\right|})}{v_{F}}\,.$ (29b)
In Eq. (28), $\bm{q}$ denotes the momentum transfer between the phonons and the electronic particle-hole excitations, $\nu$ denotes the typical frequency of the fermion, $\bm{q}_{2D}$ is the projection of $\bm{q}$ onto the $xy$-plane, $a_{z}$ is the lattice constant in the $z-$direction and ${\mathcal{N}}$ is the electron density of states on the Fermi surface. As discussed in the main text, Eq. (28) reproduces Allen’s result [31] in the limit where we drop the electron self-energy and for $c\ll v_{F}$, which allows us to drop the $A,~{}B$ terms above.
Next, let us discuss the effects of including the fermion self-energies.
(a) We first focus on the elastic scattering processes whose effects are primarily encoded in $A$. The elastic scattering rate $\Gamma$ introduces a new energy scale $T_{\rm{elastic}}=c\Gamma/v_{F}$, such that for $T<T_{\rm{elastic}}$, $\kappa$ crosses over from $T^{4}$ to $T^{5}$. However, this energy scale may not be relevant to actual experiments, because typical values of $\Gamma$ are of order $10\rm{K}$ and $c/v_{F}\sim 10^{-3}$ in cuprates. However, in systems where the Fermi velocity is small such as disordered, flat-band systems, this energy scale might be experimentally accessible, which we leave for future investigation.
(b) The second process is the inelastic scattering process of marginal Fermi liquid. Because the $\tanh(...)$ factors in Eq. (28) imply that the fermion frequency $\nu$ is comparable to phonon energy $c{\left|\bm{q}\right|}$, it has a weak effect in the $A$ factor which only leads to a small decrease of the integral at the order of $c/v_{F}$. The real part of the fermion self-energy (encoded in the factor $B$), on the other hand, can lead to a potential enhancement due to the divergence of the integrand when $B\sim{\left|\bm{q}\right|}$. Rewriting Eq. (29b) as $B\sim c{\left|\bm{q}\right|}/Z(c{\left|\bm{q}\right|})v_{F}$, where $Z(c{\left|\bm{q}\right|})\sim 1/\ln(E_{F}/c{\left|\bm{q}\right|})$ is the energy-dependent quasiparticle residue of the marginal Fermi liquid, we find that $B$ only becomes comparable to ${\left|\bm{q}\right|}$ at exponentially low-energy scales, $E_{*}\sim E_{F}\exp(-v_{F}/c)$. Therefore, for the typical range of temperatures in the normal state that are experimentally relevant between the superconducting $T_{c}$ and $O(100~{}\rm{K})$, the energy relaxation rate due to the above coupling in a MFL yields results that are nearly identical to the Fermi liquid.
I.2 Linear coupling between collective mode and phonons
We now turn to the contribution to energy relaxation due to ${\mathcal{L}}_{ep}^{(2)}$ in Eq. (4b). The corresponding phonon self-energy is
$\Pi^{\gtrless}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})=\hat{N^{i}}(\bm{x},\bm{x}^{\prime})\hat{N^{j}}(\bm{x},\bm{x}^{\prime})D_{\varphi}^{\gtrless}(t,\bm{x};t^{\prime},\bm{x}^{\prime})\,,$ (30)
where $D_{\varphi}$ is the $\varphi$-field Green’s function whose spectral function is given by Eq. (7) of the main text. Substituting this into Eq. (21), we obtain Eq. (8) of the main text.
I.3 Non-linear coupling between collective mode and phonons
The phonon self-energy is
$\Pi^{\gtrless}_{ij}(t,\bm{x};t^{\prime},\bm{x}^{\prime})=\frac{-i}{2}\hat{L^{i}}(\bm{x},\bm{x}^{\prime})\hat{L^{j}}(\bm{x},\bm{x}^{\prime})D_{\varphi}^{\gtrless}(t,\bm{x};t^{\prime},\bm{x}^{\prime})D_{\varphi}^{\lessgtr}(t^{\prime},\bm{x}^{\prime};t,\bm{x})\,.$ (31)
Compared to (22), there is a relative $-1/2$ numerical prefactor which is due to the absence of fermion loop and the symmetry factor. Following similar procedures as that for (22), we obtain Eq. (10) of the main text.