### v2

# Phonon Induced Energy Relaxation in Quantum Critical Metals

Abstract
Metals at the brink of electronic quantum phase transitions display high-temperature superconductivity, competing orders, and unconventional charge transport, revealing strong departures from conventional Fermi liquid behavior. Investigation of these fascinating intertwined phenomena has been at the center of research across a variety of correlated materials over the past many decades. A ubiquitous experimental observation is the emergence of a universal timescale that governs electrical transport and momentum relaxation. In this work, we analyze an equally important theoretical question of how the energy contained in the electronic degrees of freedom near a quantum phase transition relaxes to the environment via their coupling to acoustic phonons. Assuming that the bottleneck for energy dissipation is controlled by the coupling between electronic degrees of freedom and acoustic phonons, we present a universal theory of the temperature dependence of the energy relaxation rate in a marginal Fermi liquid. We find that the energy relaxation rate exhibits a complex set of temperature-dependent crossovers controlled by emergent energy scales in the problem. We place these results in the context of recent measurements of the energy relaxation rate via non-linear optical spectroscopy in the normal state of hole-doped cuprates.
Introduction.- In recent years, the problem of strange metals has drawn significant attention due to their unconventional transport behavior [1, 2, 3]. An experimental hallmark of these correlated metallic systems is a linear-in-temperature ($T$) electrical resistivity governed by a so-called Planckian relaxation rate of order $k_{B}T/\hbar$ down to low temperatures —in stark contrast to an umklapp-dominated $T^{2}$-scaling predicted by Fermi liquid theory, or a $T^{5}$-Bloch-Grüneissen regime due to scattering off acoustic phonons [4, 5]. The electrical transport timescale reflects the rate at which momentum is lost from the electronic system to external degrees of freedom, such as the lattice or impurities. Given the scale-invariant nature of the Planckian scattering rate [6] and experimental evidence for electronic quantum criticality across these materials [7, 8, 9, 10], a vast majority of theoretical efforts have focused on a scenario involving a Fermi surface coupled to the long-wavelength fluctuations of a bosonic collective mode of electronic origin [11, 12, 13], e.g. an electronic nematic [14]. The strong quantum critical fluctuations destroy the long-lived electronic quasiparticles near the Fermi surface, giving rise to non-Fermi liquid behavior. Early important theoretical progress [15, 16, 17] already captured many experimental signatures of the strange metals in cuprates at a phenomenological level, especially the linear-in-$T$ single-particle scattering rate. However, this does not automatically yield a $T$-linear momentum-relaxation rate, which has remained one of the central mysteries in the field. A number of theoretical works over the past few years have suggested additional microscopic ingredients that could potentially resolve this mystery [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
Figure 1: (a) The bottleneck for energy relaxation from the electronic d.o.f. at local temperature $T_{e}$ to the phonons at local temperature $T_{p}$ is governed by $\Gamma_{ep}$, which is controlled by the processes illustrated in (b)-(d). The lowest-order Feynman diagrams for the phonon self-energy due to the couplings: (b) ${\mathcal{L}}_{ep}^{(1)}$ Eq. (3), (c) ${\mathcal{L}}_{ep}^{(2)}$ Eq. (4), and (d) ${\mathcal{L}}_{ep}^{(3)}$ Eq. (5), respectively. Solid (dashed) lines denote renormalized electron (bosonic collective mode, $\varphi$) Green’s function.
In this work, we focus on a related but even more poorly understood question: How does a strongly correlated metal relax its energy to the environment? Previous studies have focused a lot on how an electron becomes incoherent through inelastic scattering, but most of the transferred energy is assumed to remain in the electronic system. In this work, we query how this energy ultimately dissipates into the environment through various symmetry-allowed couplings to phonons. While it is unlikely that phonons by themselves are responsible for the anomalous low-temperature momentum-relaxation in the strange-metal in the vicinity of quantum criticality [34], they are centrally important for energy relaxation (Fig. 1a). Phonons certainly act as a bath for the combined system consisting of electrons coupled to the bosonic collective mode, but have not received much theoretical attention in the context of non-Fermi liquids. The classic theoretical treatment [35] of energy relaxation rates, $\Gamma_{E}$, in a three-dimensional Fermi liquid coupled to acoustic phonons predicts a crossover from $\Gamma_{E}\propto T^{3}$ at low temperatures to $\Gamma_{E}\propto 1/T$ at high temperatures. In general, the momentum relaxation rate and $\Gamma_{E}$ are not automatically identical. How this behavior is modified in a non-Fermi liquid without long-lived electronic quasiparticles is presently unclear. While our results are broadly applicable to a variety of correlated electron systems in the vicinity of quantum phase transitions, it is partly inspired by the quantum critical scenario in the cuprates, which remains an active area of investigation [36, 37, 38, 39, 40, 41, 42, 43]; see also alternative points of view [44, 45, 46, 47, 48]. We note that the theoretical framework developed below can be easily adapted to correlated insulating states that host Fermi surfaces of emergent neutral fermions (e.g. spinons) coupled to dynamical gauge-fields [49, 50], assuming that these degrees of freedom (d.o.f.) also relax their energy via coupling to acoustic phonons [51].
Model of thermal relaxation.- We propose a minimal low-energy model consisting of three d.o.f.: electrons, a bosonic collective mode of electronic origin tuned to the vicinity of quantum criticality, and acoustic phonons (Fig. 1a). Our goal is to understand the temperature dependence of energy transfer from the combined electronic degrees of freedom to the phonons. The new ingredient relative to previous considerations is the presence of the bosonic collective mode, which we expect to couple to phonons via symmetry-allowed channels. The direct coupling between the bosonic collective mode and the phonons in the low-energy effective theory are associated with a distinct kinematic regime, compared to any usual local electron-phonon coupling. We investigate the effects of three distinct couplings: (i) a conventional electron–phonon deformation potential, and (ii) linear and (iii) quadratic couplings between the bosonic mode and phonons, each motivated by symmetry and microscopic considerations. We thus treat all couplings as weak perturbations compared to dominant electronic interactions.
To model the thermal relaxation process, we use the standard two-temperature model (illustrated in Fig. 1) which ignores the coupling between the phonons and the environment $\Gamma_{\text{p-env}}$, and assumes that the electrons and the phonons scattering rapidly within themselves and separately equilibrate first (with temperatures $T_{e}$ and $T_{p}$ respectively) and the relaxation bottleneck is set by the electron-phonon coupling. This is summarized by the hierarchy of scattering rates $\Gamma_{ee},\Gamma_{pp}\gg\Gamma_{ep}\gg\Gamma_{\text{p-env}}$ (see [52] for further discussion). The energy-flux between the electron and the phonon sector is given by $C_{p}{\rm d}T_{p}/{\rm d}t=-C_{e}{\rm d}T_{e}/{\rm d}t=\kappa(T_{e}-T_{p})$, and the energy relaxation rate is given by
$\Gamma_{E}=\kappa\left(\frac{1}{C_{e}}+\frac{1}{C_{p}}\right)\,$, (1)
where $C_{e(p)}$ is the specific heat of the eletrons (phonons). Note that the parametrically different dependence of $C_{e},~C_{p}$ on temperature inevitably accounts for some of the crossovers in $\Gamma_{E}$. To disentangle the thermodynamic $T$-dependent contributions to $C_{e},~C_{p}$ from the intrinsic $T$-dependence of the energy flux ($\kappa$) from electrons to phonons, in the remainder of this manuscript we focus primarily on $\kappa$ due to the distinct couplings described above. In this work, we compute $\kappa$ using the Keldysh kinetic equation [52, 53, 54], and an alternative approach based on the Kubo formula is developed in [55]. We will return to the full $T$-dependence of $\Gamma_{E}$ when we place our results in the context of the recent experiments [56].
Low-energy theory.- Consider a low-energy Lagrangian for electrons in the vicinity of quantum criticality, coupled to acoustic phonons via ${\mathcal{L}}={\mathcal{L}}_{e}+{\mathcal{L}}_{p}+{\mathcal{L}}_{ep}$, where
${\mathcal{L}}_{e}$ $=$ $\int{\rm d}\tau\Big[\sum_{{\bm{k}}}c_{{\bm{k}}}^{\dagger}\left(\partial_{\tau}+\varepsilon_{{\bm{k}}}\right)c_{{\bm{k}}}$ (2a)
$+\sum_{{\bm{q}}}\varphi_{-{\bm{q}}}\left(\partial_{\tau}^{2}+v_{\varphi}^{2}{\bm{q}}^{2}+r\right)\varphi_{{\bm{q}}}$
$+g\sum_{{\bm{k}},{\bm{q}}}f_{{\bm{k}},{\bm{{q}}}}c_{{\bm{k+q/2}}}^{\dagger}c_{{\bm{k-q/2}}}\varphi_{{\bm{q}}}\Big]+$...,\,
${\mathcal{L}}_{p}$ $=$ $\int{\rm d}\tau\sum_{i}\sum_{{\left|{\bm{q}}\right|}<{\omega_{\rm D}}/c}X_{i,-{\bm{q}}}\left(\partial_{\tau}^{2}+c^{2}{\bm{q}}^{2}\right)X_{i,{\bm{q}}}\,.$ (2b)
Here, $c_{{\bm{k}}}^{\dagger}$ creates a fermion with momentum $\bm{k}$} and dispersion $\varepsilon_{{\bm{k}}}$, and $\varphi_{{\bm{q}}}=\varphi_{-{\bm{q}}}^{\dagger}$ denotes a critical bosonic mode (e.g., Ising-nematic order [14]) near a quantum critical point (QCP) tuned by $r$. Microscopically, the $\varphi$ field arises from Hubbard-Stratonovich decoupling of the electron interaction. The boson velocity $v_{\varphi}$ is comparable to the Fermi velocity $v_{F}$, and it couples to fermions via a Yukawa coupling $g$ and form factor $f_{{\bm{k}},{\bm{q}}}$ (e.g., $B_{1g}$ irreducible representation for the Ising-nematic). $X_{i}$ represents the $i$-th component of lattice displacement, with sound speed $c$, and we adopt a Debye phonon model with cutoff $\omega_{D}/c$. For simplicity, all phonon modes are taken to have the same velocity. The $\dots$ in Eq. (2a) includes generic perturbations such as disorder.
In the clean limit, the theory yields a Landau-damped boson with dynamical exponent $z_{\varphi}=3$ and a non-Fermi liquid with $z_{c}=3/2$ [11, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 22, 89, 90, 91, 92, 20, 93, 23]. Motivated by extensive experimental evidence [13], we focus on a regime where electrons display marginal Fermi liquid (MFL) behavior [15], with self-energy $\Sigma_{c}(i\omega)\sim-\omega\ln\omega$. This can arise from various mechanisms including disorder [88, 22, 21], which reduce the boson exponent to $z_{\varphi}=2$. Importantly, our analysis of energy relaxation via phonons is agnostic to the microscopic origin of MFL, assuming only that electron–phonon couplings are weak.
Motivated by experiments in cuprates and other quasi-two-dimensional strange metals, we assume that the electronic degrees of freedom are effectively two-dimensional, with negligible interlayer hopping, while phonons remain fully three-dimensional. We consider three symmetry-allowed electron-phonon couplings, ${\mathcal{L}}_{ep}={\mathcal{L}}_{ep}^{(1)}+{\mathcal{L}}_{ep}^{(2)}+{\mathcal{L}}_{ep}^{(3)}$, where:
${\mathcal{L}}_{ep}^{(1)}$ $=$ $\int{\rm d}\tau\sum_{{\bm{k}},{\bm{q}}}M^{i}_{{\bm{k}},{\bm{q}}}c^{\dagger}_{{\bm{k+q/2}}}c_{{\bm{k-q/2}}}X_{i,{\bm{q}}}\,$, (3)
${\mathcal{L}}_{ep}^{(2)}$ $=$ $\int{\rm d}\tau\sum_{{\bm{q}}}N_{{\bm{q}}}^{i}X_{i,-{\bm{q}}}\varphi_{{\bm{q}}}\,$, (4)
${\mathcal{L}}_{ep}^{(3)}$ $=$ $\frac{1}{2}\int{\rm d}\tau\sum_{{\bm{k}},{\bm{{q}}}}L^{i}_{{\bm{k}},{\bm{q}}}\varphi_{-{\bm{k-q/2}}}\varphi_{{\bm{k+q/2}}}X_{i,{\bm{q}}}\,.$ (5)
Here, ${\mathcal{L}}_{ep}^{(1)}$ is the standard deformation-potential coupling, where $M^{i}_{{\bm{k}},{\bm{q}}}$ encodes strain coupling to the electron density. ${\mathcal{L}}_{ep}^{(2)}$ describes a linear coupling between strain and the bosonic collective mode $\varphi$, allowed by symmetry (e.g., nematic modes and $B_{1g}$ strain [94]). ${\mathcal{L}}_{ep}^{(3)}$ captures inelastic scattering between phonons and pairs of $\varphi$ bosons [95]. By Goldstone’s theorem, the matrix elements $M^{i}_{{\bm{k}},{\bm{q}}},N^{i}_{{\bm{q}}},L^{i}_{{\bm{k}},{\bm{q}}}$ scale linearly with the phonon momentum $\bm{q}$}. The theory introduced above can be formally controlled using an extension of the large-$N$ Yukawa-SYK model [3, 88, 22, 21, 96, 93, 23, 87, 33], which we review in the End Matter.
For the electron coupling $M^{i}_{{\bm{k}},{\bm{q}}}$, only the $\bm{q}$}-scaling is relevant; the details of the Fermi surface geometry are largely unimportant. In contrast, the quasi-2D nature of $\varphi$ strongly affects the boson-phonon couplings. Summing over polarizations, we model the matrix elements as $\sum_{i}|N^{i}_{{\bm{q}}}|^{2},\sum_{i}|L^{i}_{{\bm{k}},{\bm{q}}}|^{2}\sim A_{N,L}({\left|{\bm{q_{\text{2D}}}}\right|}^{2}+\lambda q_{z}^{2})$, where $A_{N}$ and $A_{L}$ have dimensions $\text{energy}]^{2}$ and $\text{energy}]^{0}$, respectively. We assume $\lambda\ll 1$ to reflect weak coupling to out-of-plane strain, and comment later on its effects.
Throughout, we assume the couplings in ${\mathcal{L}}_{ep}$ are weak compared to electronic or phononic self-interactions, justifying a perturbative treatment of the energy flux $\kappa$. While ${\mathcal{L}}_{ep}^{(2)}$ could in principle be eliminated via a change of basis for $\varphi$, we retain it explicitly, since we treat phonons as the primary channel for energy dissipation into the environment. Removing it would spuriously couple $\varphi$ directly to the bath.
The contribution $\kappa^{(1)}$ from the deformation-potential electron-phonon coupling reproduces Allen’s well-known result, i.e. a crossover from $T^{4}$ to constant near the Debye frequency $\omega_{D}$ [35]. We review this in the End Matter and also demonstrate that the MFL self-energy does not change this picture. Fig. 2 summarizes the contributions $\kappa^{(2,3)}$ from the new boson-phonon couplings proposed here, which shows a richer $T$-dependent structure that we analyze below.
Figure 2: Different regimes of the energy flux due to phonon-boson coupling (Eqs. (4),(5)) near the QCP. $\omega_{D}$ is the Debye frequency of the phonons; $c$ is the speed of sound; $v_{\varphi}$ is the velocity of the $\varphi$ boson; $\gamma$ is the Landau damping coefficient (see Eq. (6)); $m_{0}$ is the renormalized boson mass which vanishes at the QCP.
Energy-relaxation due to linear coupling between phonon and collective mode.- Let us now turn to the contribution due to ${\mathcal{L}}_{ep}^{(2)}$ in Eq. (4), with a phonon self-energy shown in Fig. 1(c). As noted previously, the renormalized propagator for $\varphi$ in the MFL is overdamped with $z_{\varphi}=2$, with a spectral function,
$A_{\varphi}(\omega,{\bm{q}})=\frac{-2\gamma\omega}{\gamma^{2}\omega^{2}+[{\left|{\bm{q_{\text{2D}}}}\right|}^{2}v_{\varphi}^{2}+m^{2}(T)]^{2}}\,.$ (6)
Here, $\gamma={\mathcal{N}}g^{2}/\Gamma$ is the Landau-damping coefficient with dimension of energy ($\mathcal{N}$}: fermion density of states, $\Gamma$: fermion elastic scattering rate), expected to be an electronic energy scale. The boson mass term, $m^{2}(T)$ includes the corrections due to a finite temperature, that are potentially important in the quantum critical regime. The QCP corresponds to $m(T=0)=0$, and the thermal fluctuations induce a $m^{2}(T)\sim\gamma T\ln(\gamma/T)$ [88, 97].
Assuming the $\varphi$ boson is equilibrated with the electrons at $T_{e}\approx T_{p}\equiv T$, the energy transfer rate takes the form $\partial_{t}E=\kappa^{(2)}(T_{e}-T_{p})$, where
$\kappa^{(2)}=\frac{1}{4}\sum_{i}\int\frac{d^{3}{\bm{q}}}{(2\pi)^{3}}|N^{i}_{{\bm{q}}}|^{2}\,[-A_{\varphi}(c|{\bm{q}}|,{\bm{q}})]\,\frac{c|{\bm{q}}|}{2T^{2}\sinh^{2}(c|{\bm{q}}|/2T)}.$ (7)
For $T\ll\omega_{D}$, we approximate $|{\bm{q}}|\sim T/c$ and, using Goldstone’s theorem, take $\sum_{i}|N^{i}_{{\bm{q}}}|^{2}\approx A_{N}{\left|{\bm{q_{\text{2D}}}}\right|}^{2}$. This yields
$\kappa^{(2)}\approx\frac{A_{N}}{c^{5}}\int_{0}^{\pi}d\theta\,\frac{\gamma T^{5}\sin^{3}\theta}{\gamma^{2}T^{2}+\left[m^{2}(T)+T^{2}v_{\varphi}^{2}\sin^{2}\theta/c^{2}\right]^{2}},$ (8)
where $\theta$ is the angle between $\bm{q}$} and the $z$-axis.
$\kappa^{(2)}$ exhibits a complex sequence of crossovers marked by different power-laws as a function of temperature, as illustrated in Fig. 2 and Fig. 3 (a). They are as follows:
(i) Let us start slightly away from the QCP, $m(T=0)=m_{0}>0$, which is assumed to be smaller than all other energy scales, including $\omega_{\rm D}$}. At the lowest temperature, the denominator of Eq. (8) is dominated by $m_{0}$, so we obtain $\kappa^{(2)}\sim(A_{N}/c^{5})(\gamma/m_{0}^{4})T^{5}$. In this regime, $\varphi$ mediates local interaction between the electrons, and the effects of damping and dispersion are suppressed relative to $m_{0}$. Alternatively, the result can be derived from $\kappa^{(2)}\sim A_{N}T^{4}\times\braket{\varphi\varphi}$, where the $T^{4}$ prefactor is the result of phonon phase counting and Gold-stone coupling. In the low-energy and long-wavelength limit, the boson $\varphi$ is expected to behave as an ohmic bath, which yields $\braket{\varphi\varphi}\propto\gamma T/m_{0}^{4}$. We also remark in passing that similar bosonic collective modes are generically present in interacting Fermi liquids, and $\kappa^{(2)}$ can be interpreted as an interaction correction to energy relaxation on top of the more dominant free fermion result $\kappa^{(1)}$. When the system is deep in the Fermi liquid phase, the boson mass $m_{0}$ is expected to be an electronic energy scale that overwhelms other crossovers, i.e. $\kappa^{(2)}$ grows as $T^{5}$ until being cutoff at $T\sim{\omega_{\rm D}}$.
(ii) With increasing $T>m_{0}^{2}/\gamma$, the $\gamma^{2}T^{2}$ term in the denominator of Eq. (8) dominates, so $\kappa^{(2)}\sim(A_{N}/c^{5})T^{3}/\gamma$, which becomes comparable to $\kappa^{(1)}$, albeit with a different coefficient. In this regime, the $z_{\varphi}=2$ dynamics first rears its head; however, the typical phonon momentum is small compared to the typical $\varphi$ frequencies. Therefore, the phonons are more sensitive to the homogeneous fluctuations of $\varphi$, as opposed to their full diffusive character. Mathematically, regimes (i) and (ii) can be described in a unified fashion by defining $\Delta(T)=\max(m(T),\sqrt{\gamma T})$ and writing $\kappa^{(2)}\sim(A_{N}/c^{5})\gamma T(T/\Delta(T))^{4}$; this applies to the quantum critical fan ($m_{0}=0$) as well as for $\Delta(T)\to m(T)$.
(iii) When $T\gtrsim\gamma(c^{2}/v_{\varphi}^{2})$, the $T^{2}v_{\varphi}^{2}/c^{2}$ term in the denominator becomes relevant. The $\theta$-integral in Eq. (8) should be cutoff at $\theta_{\text{min}}\sim(c/v_{\varphi})\Delta(T)/T$, which yields $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma T\ln(v_{\varphi}^{2}T/(c^{2}\gamma))$ (see End Matter), which signals an enhanced contribution due to phonons moving in the $z$ direction. In this regime, the typical phonon momentum is large compared to the typical $\varphi$ frequencies, and unlike the previous lower-temperature regime, the phonons are now sensitive to diffusive character of $\varphi$. Such a regime is unique to phonons: for a bosonic mode of electronic origin with $c\sim v_{\varphi}$, it would only emerge at $T\sim\gamma$, an electronic scale. Here, it appears at much lower temperatures due to $c\ll v_{\varphi}$.
(iv) With a further increase $T\gtrsim{\omega_{\rm D}}$, the integral saturates, and we obtain a constant $\kappa^{(2)}\sim A_{N}/(cv_{\varphi}^{4})\gamma\omega_{D}\ln(v_{\varphi}^{2}\omega_{D}/(c^{2}\gamma))$.
It is useful to make a few remarks about the underlying dimensionality of the problem. The lowest temperature regimes (i), (ii) exhibiting $\kappa^{(2)}\sim T^{5}$ and $\kappa^{(2)}\sim T^{3}$ depend on the phonon dimensionality in a straightforward fashion via their density of states. However, the $\kappa^{(2)}\sim T\ln(T)$ regime (iii) involves a subtle kinematic interplay between the phonon and $\varphi$. In particular, the phonon dispersion along the $z$-direction contributes more significantly compared to the other directions, leading to the $\ln$ enhancement; the latter would be absent if the phonons were purely two-dimensional. Furthermore, $\kappa^{(2)}$ in this regime also depends crucially on the assumption of completely decoupled nature of the two-dimensional electronic layers, via the simplification, $N^{i}_{{\bm{q}}}\propto{\left|{\bm{q_{\text{2D}}}}\right|}$, instead of $\left|{\bm{q}}\right|$}. Finally, let us also consider the additional perturbation due to coupling the $\varphi$ boson to out-of-plane strain, which leads to a correction, $\lambda q_{z}^{2}$ term in $|N^{i}_{{\bm{q}}}|^{2}$ with $\lambda\ll 1$. This does not lead to qualitatively different results in regimes (i) and (ii), but it can enhance the $\ln$ factor in regime (iii) to a linear factor of $v_{\varphi}^{2}T/(c^{2}\gamma)$ (see End Matter). Since this regime is upper bounded by $T\sim\omega_{D}$, this term remains a small perturbation if $\lambda\ll\gamma c^{2}/(\omega_{D}v_{\varphi}^{2})$.
Figure 3: Numerical plot of (a) $\kappa^{(2)}(T)$ and (b) $\kappa^{(3)}(T)$ on log-log scale. The blue curves are numerical evaluation of Eqs. (7) and (9). The dashed lines denote the approximate scaling forms deduced analytically.
Energy-relaxation due to non-linear coupling between phonon and collective mode.- Finally, we consider the contribution due to ${\mathcal{L}}_{ep}^{(3)}$ in Eq. (5), with a phonon self-energy shown in Fig. 1(d), which involves scattering between $\varphi^{2}$ and the acoustic phonons. The general expression for the flux $\kappa^{(3)}$ reads
$\begin{split}&\kappa^{(3)}=-\frac{1}{16}\int_{{\bm{q}}}\frac{{\rm d}^{3}{\bm{q}}}{(2\pi)^{3}}\frac{c{\left|{\bm{q}}\right|}}{2T^{2}\sinh^{2}\frac{c{\left|{\bm{q}}\right|}}{2T}}\int\frac{{\rm d}\nu}{2\pi}\int\frac{{\rm d}^{3}{\bm{k}}}{(2\pi)^{3}}\\ &\times\sum_{i}|L^{i}_{{\bm{k}}+{\bm{q}}/2,{\bm{q}}}|^{2}A_{\varphi}(\nu,{\bm{k}})A_{\varphi}(\nu+c{\left|{\bm{q}}\right|},{\bm{k}}+{\bm{q}})\\ &\times\left[\coth\bigg(\frac{\nu+c{\left|{\bm{q}}\right|}}{2T}\bigg)-\coth\bigg(\frac{\nu}{2T}\bigg)\right]\,.\end{split}$ (9)
Once again, there is a complex sequence of temperature dependent crossovers that can be obtained with the following approximations: replace $\sum_{i}|L^{i}_{{\bm{k}}+{\bm{q}}/2,{\bm{q}}}|^{2}$ by $A_{L}{\left|{\bm{q_{\text{2D}}}}\right|}^{2}$, and for $T\lesssim{\omega_{\rm D}}$, include contributions from phase-space where $c{\left|{\bm{q}}\right|}\sim\nu\sim T$. We obtain
[…]
where the vector ${\bm{q}}_{\text{2D}}(\theta)$ has norm $(T/c)\sin\theta$, and $\theta$ is the angle between $\bm{q}$} and the z-axis.
[…]
Connection to Experiments.- Recent optical experiments involving the technique of two-dimensional coherent spectroscopy [98, 99, 100] have studied the nonlinear response of correlated metals and its relationship to energy relaxation [101, 56], providing a further impetus to a careful theoretical examination of the questions raised above. Over a broad range of hole-dopings in a specific cuprate family [102, 103], recent pump-probe measurements [56] have identified energy relaxation as the dominant nonlinear process that governs the experimental phenomenology. At low temperatures across most of the doping phase diagram, $\Gamma_{E}\sim k_{B}T/\hbar$, with a rate that is significantly slower (approximately by an order of magnitude) compared to the Planckian momentum-relaxation rate [1, 104]. With increasing temperature, there is a tendency for $\Gamma_{E}$ to saturate across a doping dependent crossover scale. Both of these regimes are in stark contrast to the results expected in a conventional Fermi liquid coupled to acoustic phonons [35]. In THz experiments, it is the second rate $\Gamma_{ep}$ that is measured, which leads to a fast initial decay of the pump-probe signal to a plateau at the order of pico seconds, whereas the phonon-environment relaxation $\Gamma_{\text{p-env}}$ leads to a slower decay on longer timescales.
To make contact with experiments, we convert $\kappa=[\kappa^{(1)}+\kappa^{(2)}+\kappa^{(3)}]$ to the energy relaxation rate, $\Gamma_{E}$, via Eq. (1). Here, due to the comparison between the heat capacities $C_{p}\sim(T/{\omega_{\rm D}})^{3}$ and $C_{e}\sim(T/E_{F})\ln(\gamma/T)$, an additional crossover is present, whose temperature scale $T_{*}\sim\sqrt{{\omega_{\rm D}}^{2}/E_{F}}$ is set by equating $C_{p}$ and $C_{e}$. Eq. (1) states that the sector with the smaller heat capacity dominates as its temperature changes more rapidly. When $T\ll T_{*}$, we have $\Gamma_{E}\approx\kappa/C_{p}$ and when $T\gg T_{*}$, $\Gamma_{E}\approx\kappa/C_{e}$. From thermodynamic measurements in a representative family of the cuprates [37], $T_{*}$ is at the order of 11-1010 K. For the experimental range of $T\sim O(100\textrm{ K})$, the latter is more relevant, so we would then divide the scaling of $\kappa$ by a power of $T$ to get scaling of $\Gamma_{E}$. Our theoretical results for $\kappa^{(3)}$ provide an appealing scenario for the experimental observations. In the experimental temperature range, the energy relaxation rate $\Gamma_{E}$ is related to $\kappa^{(3)}$ by $\Gamma_{E}\approx\kappa^{(3)}/C_{e}$. Therefore, in regime (iii) and regime (iv) we would obtain a crossover from $T\ln T$ to $T$-independent constant in $\Gamma_{E}$, which qualitatively agrees with the observation in [56] for samples near optimal doping (see Fig. 6c in [56]).
Discussion and Outlook.- We have shown that near an electronic QCP, symmetry-allowed couplings to acoustic phonons generate rich, temperature-dependent crossovers in the energy relaxation rate. When momentum relaxation in a marginal Fermi liquid is governed by electronic processes and disorder, while energy relaxation is limited by phonon coupling, there is no a priori reason for the two relaxation rates to exhibit a similar behavior. Our analysis reveals that energy relaxation follows distinct scaling behaviors across emergent low-energy regimes, independent of momentum relaxation. We also note that (a) while the electron-phonon coupling, can in principle, modify the structure of the underlying QCP [94, 95, 105, 106], the interesting regime (iii) obtained above will not be affected, and (b) the energy relaxation process we address is not affected by potential boson drag effects [107, 108, 20, 22]; see [52] for details.
Recent experiments [109, 110, 111] have motivated proposals that phonons in strongly correlated metals may lose their quasiparticle character, exhibiting Planckian lifetimes $O(\hbar/k_{B}T)$ [107, 112, 113, 114, 108]. Despite this broadening, our results for energy flux should still apply, as the relevant phonon frequencies remain $O(k_{B}T)$. A full theoretical treatment of a non-Fermi liquid entangled with incoherent phonons is left for future work. Additionally, it would be interesting to investigate how the phonon dynamics is modified by the non-Fermi liquid through numerical approaches, such as quantum Monte-Carlo [31, 12, 80, 115, 116].
In this work we have focused on the normal state, but pairing fluctuations above $T_{c}$ and Bogoliubov quasiparticles below $T_{c}$ may also influence energy relaxation through their coupling to phonons [117], an interesting direction for future study. Moreover, the complex crossovers we identify—arising from mismatched velocities and dimensionality of phonons and collective modes—suggest similar behavior may appear in sound attenuation near quantum criticality [118, 119, 120, 121, 122, 123, 51, 124], which we leave for future investigation.
Acknowledgments.- We thank N.P. Armitage, D. Chaudhuri, and B. Ramshaw for stimulating discussions. H.G. is supported by a Wilkins postdoctoral fellowship at Cornell University. D. C. is supported in part by a NSF CAREER grant (DMR-2237522), and a Sloan Research Fellowship.